These were approved.
Reminder: CEG has decreed that systems should be fixed promptly or removed from monitoring.
Neil did email to COs and then threw the switch. Apache will spot now-invalid configurations and so not try to restart using them. It'll be in this afternoon's <stable> release.
Carol clarified that her action was to obtain a device for physically rendering discs useless. The action to set up a zeroing station was really on Iain.
The question of where to put it was raised, but not resolved.
Most are done. Amend the action to refer only to git and gerrit.
Re-defer to August.
Still a few to do.
It would make sense to defer this to after the "IPv6 for student labs" action.
Actions added or revived:
Report from Computing Executive Group
This is mandatory.
Reports from Units
You can see examples of this by following the links in the report.
at9camsystem (RT#88987). Moving the volume to a faster server has helped a bit. It looks as though it might be a resourcing issue. Still under investigation.
We have one student helper lined up for the AT-6.06 work. More would be good to have!
Topics for discussion
Over the years we have had cleaning staff and contractors working unsupervised in the Forum server room, which we have been reasonably happy with to date, and separate contractors to empty B.03 periodically. The latter group definitely do need to be supervised to ensure that they are interpreting their instructions in the way we expect.
The issue must arise elsewhere. George will ask Paul Hutton what happens in IS server rooms, such as the three AT basement rooms. Meanwhile, use common sense.
This prompted much discussion, covering inter alia
There was a suggestion that something to say why we have the versions we do, and how to get hold of newer ones, might be useful. Take to CEG to discuss further and finger someone.
Q: is our Data Protection and Interception Statement on the computing.help site meant to be current and definitive, and something that can be linked to?
A: yes. (It is still in need of a GDPR overhaul though.)
Specifically, there was an RT ticket asking for contact details in an unmaintained personal web page to be updated. Craig will take the issue of unmaintained pages to the Web Strategy Group.
The next meeting will be on Wednesday 23rd May in IF-4.31, at 10:00.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh